OpenStudy (mathmath333):

a new flag is to be designed with six verticle stripes using some of all the colors yellow,green , blue and red,then the number of ways this can be done such that no two adjacent stripes have the same color is a.12* 81 b.16*192 c.20*125 d.24*216

OpenStudy (anonymous):

For the first stripe, you have 4 choices. For the second, you have 3 choices, since the previous color cannot be picked. For the third, again you have 3 choices, since it cannot have the same color as the second stripe, but it can take on the color of the first. The remaining stripes likewise have only 3 choices.

OpenStudy (mathmath333):

so u mean \(\large\tt \begin{align} \color{black}{4\times3\times 3\times 3\times 3\times 3\\ =4\times 3^5\\ =12\times 81 }\end{align}\) can u confirm this ?

OpenStudy (anonymous):

That's right. Not sure why the possible choices are listed like that.

OpenStudy (mathmath333):

Ah ok thanks , this for tough entrance exam so they might confuse students